OpenStudy (anonymous):

4. Gavin and Seiji both worked hard over the summer. Together, they earned a total of $425. Gavin earned $25 more than Seiji. How much did each of them earn?

OpenStudy (anonymous):

x = Gavin y = Seiji x + y = 425 Gavin earned $25 more than Seiji x = (y + 25) (y + 25) + y = 425 Solve the equation for y, which is Seiji.

OpenStudy (anonymous):

What did you get for y (Seiji) ?

OpenStudy (anonymous):

Once you get the answer for y (Seiji), put the answer for y (Seiji) into this equation x = Gavin x = y + 25